Use 9.81m/s2 for gravity if needed!

1) The magnitude of the restoring force that acts on a 1.0-kg pendulum when it is displaced 2.0° from equilibrium is ___ N. 


2) The magnitude of the restoring force that acts on a 1.0-kg pendulum when it is displaced 10.0° from equilibrium is ___ N. 


3)  On earth the period of a 1.50-m pendulum is ___s.


4) On the hypothetical planet Xeon, a pendulum with a length of 95.0 cm swings with a frequency of 1.50 Hz. The acceleration due to gravity on Xeon is ___ m/s2.


5) A pendulum is swinging back and forth:


Which of the following factors affect the period of the pendulum:

I. length

II. mass

III. angle of release

IV. accleration due to gravity


A- I only


B- I and II


C- III and IV


D- I and IV





1
Expert's answer
2021-07-19T09:48:24-0400

1)


"F=(1)(9.8)\sin{2}=0.34\ N"

2)


"F=(1)(9.8)\sin{10}=1.7\ N"

3)


"T=2\pi\sqrt{\frac{1.5}{9.8}}=6.04\ s"

4)


"\frac{1}{1.5}=2\pi\sqrt{\frac{0.95}{g}}\\g=84.4\frac{m}{s^2}"

5) D- I and IV